# Lung-DDPM+: Efficient Thoracic CT Image Synthesis using Diffusion Probabilistic Model

### v1

@@ 4.4.2 Lung nodule segmentation evaluation
    Training Set Augmentation – Synthetic CT images are combined with the cross-validation training sets at varying synthetic-to-real ratios (100% to 1000%) to create a mixed training set for training a U-Net [31] segmentation model.
@@
Lung-DDPM [8] +100% 0.5083±0.0501 45.7542±3.2250 0.2625±0.0617 174.8173±49.8686
@@
Table 3 presents the lung nodule segmentation evaluation results, comparing traditional data augmentation methods, Lung-DDPM+, Lung-DDPM [8], GEM-3D [9], and Med-DDPM [6] across 2D and 3D segmentation tasks. The results demonstrate that Lung-DDPM+ achieves comparable segmentation performance to other SOTA models while maintaining superior computational efficiency. In 3D lung nodule segmentation, Med-DDPM achieves the highest Dice score (0.5640 at 1000% synthetic data) and the lowest HD95 (26.8428 at 1000%), while Lung-DDPM+ closely follows with a Dice score of 0.5626 at 900% and an HD95 of 24.6928, which is lower than Med-DDPM’s best HD95 score, suggesting that Lung-DDPM+ maintains strong anatomical precision. In 2D segmentation, Lung-DDPM (0.4670 at 500%) and Lung-DDPM+ (0.4460 at 500%) achieve comparable Dice scores, while Med-DDPM (HD95 = 107.4805 at 900%) maintains the lowest boundary error. Notably, Lung-DDPM+ consistently outperforms GEM-3D, which struggles to generate meaningful segmentation results, with significantly lower Dice scores (max 0.4720 at 200%) and higher HD95 values across all settings. These findings highlight that Lung-DDPM+ matches or exceeds the performance of other generative models while achieving better sampling efficiency, requiring less computational overhead than Lung-DDPM and Med-DDPM while still producing high-quality lung nodules with clear anatomical structures. Furthermore, Lung-DDPM+ effectively leverages synthetic data, demonstrating that a balanced synthetic-to-real data ratio (500%–900%) significantly boosts segmentation performance, particularly in 3D segmentation tasks, where Lung-DDPM+ exhibits a clear advantage in maintaining detailed anatomical structures. These results confirm that Lung-DDPM+ provides an optimal balance between efficiency and segmentation accuracy, making it a compelling choice for medical imaging applications that require both high-quality synthetic data generation and real-time feasibility.
@@ 4.5 Qualitative results
The Visual Turing Test was conducted on 61 randomly selected real-synthetic case pairs, comprising 61 real CT scans and 61 synthetic CT scans. Each pair included a real case, the corresponding lung nodule segmentation, and synthetic cases generated based on the segmentation. The radiologist was asked to identify which case was fake in each pair. Following the methodologies outlined in [21] and [22], all pairs were provided as 2D animations to effectively represent the 3D CT volume while focusing exclusively on the nodule-positive slices. The evaluation results indicate that the radiologist achieved an overall accuracy of 0.6393, with a sensitivity of 0.7647 and a specificity of 0.4815. These results demonstrate that while the radiologist exhibited a relatively high ability to correctly identify synthetic cases when a fake was present (high sensitivity), the specificity was notably lower, indicating challenges in distinguishing real cases when presented with the synthetic alternatives. The lower specificity suggests that certain characteristics of the synthetic images closely mimicked real CT scans, potentially leading to false-positive identifications. This discrepancy between sensitivity and specificity underscores the realism of the synthetic data, particularly in replicating the nuanced features of real lung nodules, and highlights the potential of the proposed Lung-DDPM+ in contributing to robust synthetic datasets for different downstream thoracic imaging applications.
@@ 5 Discussion
The efficiency of generative models has been widely studied and recognized as a critical factor for practical deployment [32]. However, improving efficiency remains challenging, particularly for emerging approaches with complex architectures and diverse applications. Various efforts have been made to enhance computational efficiency across SOTA generative models, including generative adversarial networks (GANs) [33], variational autoencoders (VAEs) [34], and diffusion models [35]. Despite these advancements, most efficiency improvements have been tailored to general visual tasks, rather than medical imaging applications. In this paper, we introduce Lung-DDPM+, a novel diffusion-based model for efficient lung nodule CT image synthesis, achieving significant efficiency gains over SOTA generative models. As shown in Table 2, Lung-DDPM+ drastically reduces computational overhead, requiring up to 8× fewer TFLOPs, 21× less VRAM, and 76× shorter sampling time compared to Lung-DDPM and Med-DDPM. Given its real-time generation capability, Lung-DDPM+ is particularly well-suited for cost-sensitive and edge-computing scenarios, enabling the broader adoption of high-quality synthetic medical imaging data.
@@
In addition to efficiency, anatomical precision is a crucial aspect of synthetic medical imaging. Our previous work, Lung-DDPM [8], introduced a semantic layout-guided diffusion model to generate pulmonary regions, which were then blended with extra-pulmonary areas using an anatomically aware sampling process. While Lung-DDPM partially addressed the anatomical imprecision issue, it still struggled to reconstruct fine pulmonary structures, such as vasculature and the bronchial tree, due to their small size and low contrast. Instead of synthesizing an entire pulmonary region, Lung-DDPM+ exclusively generates nodule regions, which are seamlessly integrated into surrounding tissues. This novel approach allows the model to focus on nodule texture while preserving anatomical accuracy. As demonstrated in Figure 5, this targeted synthesis strategy significantly improves structural consistency and generation stability compared to other SOTA methods.
@@
Beyond efficiency and anatomical accuracy, synthetic lung nodule CT images play a vital role in downstream applications, such as lung cancer risk prediction [30], tumor segmentation [36], and lung cancer survival prediction [37]. To assess synthetic sample quality, we evaluate Lung-DDPM+ on 2D and 3D lung nodule segmentation tasks, using Dice scores and HD95 metrics as performance indicators. As shown in Table 3, synthetic samples enhance segmentation model performance across both 2D and 3D tasks. However, achieving optimal performance requires a significant proportion of synthetic data—typically above 500%—and each model exhibits a distinct sweet spot for the ideal synthetic-to-real data ratio. Notably, 3D segmentation models demand more synthetic data than 2D models to attain meaningful improvements, emphasizing the need for carefully designed data augmentation strategies. Additionally, we observe a marginal effect, where the performance gains saturate as more synthetic data is introduced. This underscores the importance of balancing synthetic data volume, generation cost, and existing dataset scale to maximize segmentation accuracy while maintaining computational efficiency.
@@ 6 Conclusion and future study
In this paper, we propose Lung-DDPM+, a novel CT image synthesis method capable of generating high-quality nodule-positive CT images efficiently. The proposed pulmonary anatomically aware sampling process enables Lung-DDPM+ to synthesize realistic lung nodules from healthy CT slices while reducing computational complexity and hardware requirements compared to its predecessor. Additionally, Lung-DDPM+ preserves detailed and anatomically accurate structures throughout the sampling process. Experimental results demonstrate that the proposed method outperforms SOTA competitors in sampling efficiency and downstream lung nodule segmentation tasks, highlighting its potential in various medical imaging applications. In future studies, we will explore how to extend this method to a broader range of tumors, organs, and medical imaging modalities to enhance the deployability of generative models in clinical and research settings.

### v2

@@ 4.4.2 Lung nodule segmentation evaluation
    Training Set Augmentation – Synthetic CT images are combined with the cross-validation training sets at varying synthetic-to-real ratios (100% to 1000%) to create a mixed training set for training a U-Net [ronneberger2015u] segmentation model.
@@
Lung-DDPM [jiang2025lungddpm] +100% 0.5083±0.0501 45.7542±3.2250 0.2625±0.0617 174.8173±49.8686
@@
Table 3 presents the lung nodule segmentation evaluation results, comparing traditional data augmentation methods, Lung-DDPM+, Lung-DDPM [jiang2025lungddpm], GEM-3D [zhu2024generative], and Med-DDPM [dorjsembe2024conditional] across 2D and 3D segmentation tasks. The results demonstrate that Lung-DDPM+ achieves comparable segmentation performance to other SOTA models while maintaining superior computational efficiency. In 3D lung nodule segmentation, Med-DDPM achieves the highest Dice score (0.5640 at 1000% synthetic data) and the lowest HD95 (26.8428 at 1000%), while Lung-DDPM+ closely follows with a Dice score of 0.5626 at 900% and an HD95 of 24.6928, which is lower than Med-DDPM’s best HD95 score, suggesting that Lung-DDPM+ maintains strong anatomical precision. In 2D segmentation, Lung-DDPM (0.4670 at 500%) and Lung-DDPM+ (0.4460 at 500%) achieve comparable Dice scores, while Med-DDPM (HD95 = 107.4805 at 900%) maintains the lowest boundary error. Notably, Lung-DDPM+ consistently outperforms GEM-3D, which struggles to generate meaningful segmentation results, with significantly lower Dice scores (max 0.4720 at 200%) and higher HD95 values across all settings. These findings highlight that Lung-DDPM+ matches or exceeds the performance of other generative models while achieving better sampling efficiency, requiring less computational overhead than Lung-DDPM and Med-DDPM while still producing high-quality lung nodules with clear anatomical structures. Furthermore, Lung-DDPM+ effectively leverages synthetic data, demonstrating that a balanced synthetic-to-real data ratio (500%–900%) significantly boosts segmentation performance, particularly in 3D segmentation tasks, where Lung-DDPM+ exhibits a clear advantage in maintaining detailed anatomical structures. These results confirm that Lung-DDPM+ provides an optimal balance between efficiency and segmentation accuracy, making it a compelling choice for medical imaging applications that require both high-quality synthetic data generation and real-time feasibility.
@@ 4.5 Qualitative results
The Visual Turing Test was conducted on 61 randomly selected real-synthetic case pairs, comprising 61 real CT scans and 61 synthetic CT scans. Each pair included a real case, the corresponding lung nodule segmentation, and synthetic cases generated based on the segmentation. The radiologist was asked to identify which case was fake in each pair. Following the methodologies outlined in [chen2024towards] and [hu2023label], all pairs were provided as 2D animations to effectively represent the 3D CT volume while focusing exclusively on the nodule-positive slices. The evaluation results indicate that the radiologist achieved an overall accuracy of 0.6393, with a sensitivity of 0.7647 and a specificity of 0.4815. These results demonstrate that while the radiologist exhibited a relatively high ability to correctly identify synthetic cases when a fake was present (high sensitivity), the specificity was notably lower, indicating challenges in distinguishing real cases when presented with the synthetic alternatives. The lower specificity suggests that certain characteristics of the synthetic images closely mimicked real CT scans, potentially leading to false-positive identifications. This discrepancy between sensitivity and specificity underscores the realism of the synthetic data, particularly in replicating the nuanced features of real lung nodules, and highlights the potential of the proposed Lung-DDPM+ in contributing to robust synthetic datasets for different downstream thoracic imaging applications.
@@ 5 Discussion
The efficiency of generative models has been widely studied and recognized as a critical factor for practical deployment [liu2024deep]. However, improving efficiency remains challenging, particularly for emerging approaches with complex architectures and diverse applications. Various efforts have been made to enhance computational efficiency across SOTA generative models, including generative adversarial networks (GANs) [tao2022df], variational autoencoders (VAEs) [rybkin2021simple], and diffusion models [ma2024efficient]. Despite these advancements, most efficiency improvements have been tailored to general visual tasks, rather than medical imaging applications. In this paper, we introduce Lung-DDPM+, a novel diffusion-based model for efficient lung nodule CT image synthesis, achieving significant efficiency gains over SOTA generative models. As shown in Table 2, Lung-DDPM+ drastically reduces computational overhead, requiring up to 8× fewer TFLOPs, 21× less VRAM, and 76× shorter sampling time compared to Lung-DDPM and Med-DDPM. Given its real-time generation capability, Lung-DDPM+ is particularly well-suited for cost-sensitive and edge-computing scenarios, enabling the broader adoption of high-quality synthetic medical imaging data.
@@
In addition to efficiency, anatomical precision is a crucial aspect of synthetic medical imaging. Our previous work, Lung-DDPM [jiang2025lungddpm], introduced a semantic layout-guided diffusion model to generate pulmonary regions, which were then blended with extra-pulmonary areas using an anatomically aware sampling process. While Lung-DDPM partially addressed the anatomical imprecision issue, it still struggled to reconstruct fine pulmonary structures, such as vasculature and the bronchial tree, due to their small size and low contrast. Instead of synthesizing an entire pulmonary region, Lung-DDPM+ exclusively generates nodule regions, which are seamlessly integrated into surrounding tissues. This novel approach allows the model to focus on nodule texture while preserving anatomical accuracy. As demonstrated in Figure 5, this targeted synthesis strategy significantly improves structural consistency and generation stability compared to other SOTA methods.
@@
Beyond efficiency and anatomical accuracy, synthetic lung nodule CT images play a vital role in downstream applications, such as lung cancer risk prediction [jiang2025benchmark], tumor segmentation [jiang2018tumor], and lung cancer survival prediction [lai2020overall]. To assess synthetic sample quality, we evaluate Lung-DDPM+ on 2D and 3D lung nodule segmentation tasks, using Dice scores and HD95 metrics as performance indicators. As shown in Table 3, synthetic samples enhance segmentation model performance across both 2D and 3D tasks. However, achieving optimal performance requires a significant proportion of synthetic data—typically above 500%—and each model exhibits a distinct sweet spot for the ideal synthetic-to-real data ratio. Notably, 3D segmentation models demand more synthetic data than 2D models to attain meaningful improvements, emphasizing the need for carefully designed data augmentation strategies. Additionally, we observe a marginal effect, where the performance gains saturate as more synthetic data is introduced. This underscores the importance of balancing synthetic data volume, generation cost, and existing dataset scale to maximize segmentation accuracy while maintaining computational efficiency.
While our experiments were conducted using the standard U-Net architecture to maintain methodological clarity and enable controlled comparisons across generative models, we acknowledge that more advanced segmentation frameworks have emerged in recent years. Architectures such as nnU-Net [isensee2021nnu], which incorporates automated pipeline optimization and extensive built-in augmentation strategies, represent the current state-of-the-art in medical image segmentation. Future work could extend our augmentation strategy to such advanced and automated segmentation pipelines to evaluate whether the benefits of synthetic data generation persist or even amplify when combined with more sophisticated architectures. Given that nnU-Net already includes comprehensive traditional augmentation methods, investigating how model-based synthetic data augmentation complements or enhances these existing strategies would be particularly valuable. Such studies would help establish whether our approach provides additive benefits in highly optimized segmentation frameworks and could further validate the generalizability of Lung-DDPM+ across diverse clinical applications and architectural choices.
@@ 6 Conclusion and future study
In this paper, we propose Lung-DDPM+, a novel CT image synthesis method capable of generating high-quality nodule-positive CT images efficiently. The proposed pulmonary anatomically aware sampling process enables Lung-DDPM+ to synthesize realistic lung nodules from healthy CT slices while reducing computational complexity and hardware requirements compared to its predecessor. Additionally, Lung-DDPM+ preserves detailed and anatomically accurate structures throughout the sampling process. Experimental results demonstrate that the proposed method outperforms SOTA competitors in sampling efficiency and downstream lung nodule segmentation tasks, highlighting its potential in various medical imaging applications.
While our experiments employed the standard U-Net architecture for methodological clarity and controlled comparison, future work could extend our augmentation strategy to more advanced segmentation frameworks such as nnU-Net. Investigating how model-based synthetic data complements the sophisticated built-in augmentation strategies of such pipelines would help establish the broader applicability of Lung-DDPM+ across diverse architectural choices and clinical applications.
Although our nodule semantic layout generation uses ellipsoidal approximations, our results demonstrate that this simplified approach achieves substantial segmentation improvements, suggesting the diffusion model generates diverse nodule textures beyond initial geometric constraints. Future work could investigate whether incorporating more sophisticated morphological models for spiculated, ground-glass, or lobulated presentations would further enhance synthetic data quality and downstream performance.
Additionally, we will explore how to extend this method to a broader range of tumors, organs, and medical imaging modalities to enhance the deployability of generative models in clinical and research settings.